OpenStudy (anonymous):

I have an integration problem that I would like solved. I'll specify it using the equation editor in a minute. The equation editor doesn't seem to appear until after I do a post. Okay, updating this - I have the equation editor now. \[F(x) = \int\limits_{2}^{x^2}(t^{1/3})\sin(t^{2}-3t)\]

OpenStudy (anonymous):

Refresh the page if the equation editor isn't showing up.

OpenStudy (anonymous):

Okay, here is the equation. I don't know why ... but I can't see the equation editor in the original problem post box. I can only see it as an option in the reply box. So here is the integral: \[F(x) = \int\limits_{{2}}^{x^2}(t ^{\frac{ 1 }{ 3 }})\sin(t ^{2}-3t)dt\] I've tried integration by parts (but it becomes more and more complex instead of simplifying) and I've tried by substitution but I can't solve it. Perhaps there was a mistake in the original exam paper or perhaps there is a solution method/technique that I just haven't spotted yet.

OpenStudy (loser66):

Mine is ugly one but there is an outlet. \[u = t^{\frac{4}{3}} \rightarrow t ^2 = u^{\frac{3}{2}} \\and ~~ du= \frac{4}{3}t^{\frac{1}{3}}dt. \]

OpenStudy (loser66):

then after applying , you have to use formula sin(a-b)= sin a cos b - sinb cos a, to solve the problem. Good luck

OpenStudy (loser66):

don't forget convert t = u^ ( 3/4) to replace inside sin (t^2 -3t)

OpenStudy (anonymous):

@Aragusha Are you sure you are not supposed to find the derivative?

OpenStudy (amistre64):

just wondering about an integration ...... i agree that htis is prolly a derivative of F(x) problem tho \[sin(u)=\sum_0\frac{(-1)^{n}}{(2n+1)!}u^{2n+1}\] \[sin(t^2-3t)=\sum_0\frac{(-1)^{n}}{(2n+1)!}(t^2-3t)^{2n+1}\] \[t^{(1/3)~}sin(t^2-3t)=\sum_0\frac{(-1)^{n}}{(2n+1)!}(t^2-3t)^{2n+4/3}\] \[\sum_0\frac{(-1)^{n}}{(2n+1)!}(t^2-3t)^{2(n+2/3)}\]

OpenStudy (amistre64):

even the wolf is having issues with it :)

OpenStudy (anonymous):

Hi guys, thank you all for your input. The question states ..... state the Fundamental Theorem of calculus and use it to DIFFERENTIATE the following function w.r.t. x ....(the equation being the one listed above). But does this not still mean I have to integrate the right hand side?

OpenStudy (amistre64):

it does not

OpenStudy (amistre64):

\[\int_{a(x)}^{b(x)}f(t)~dt=F(b(x))-F(a(x))\] \[\frac d{dx} \int_{a(x)}^{b(x)}f(t)~dt=\frac d{dx}F(b(x))-F(a(x))\]

OpenStudy (amistre64):

using the chain rule we get back to \[=\frac d{dx}F(b(x))-F(a(x))=f(b(x))b'(x)-f(a(x))a'(x)\]

OpenStudy (amistre64):

notice that regardless of the integration, we still ahve all the parts we need to define the end results

OpenStudy (amistre64):

\[\frac{d}{dx}F(x) = \frac{d}{dx}\int\limits_{2}^{x^2}(t^{1/3})\sin(t^{2}-3t)\to\\ ~~~~~~~~~([x^2]^{1/3})\sin([x^2]^{2}-3(x^2))*2x-([2]^{1/3})\sin([2]^{2}-3(2))*2'\] \[2x(x^{2/3})\sin(x^4-3x^2)\]

OpenStudy (anonymous):

Hi Amistre, thanks for this. I didn't understand the question properly but your solution makes sense to me. It's easier than I thought ... once you know how ... which I didn't. I really appreciate your input and that of others and apologies to all who spent time on this for posing the question in a misleading way originally.

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

Thanks.